# The Artificial Benchmark for Community Detection with Outliers and Overlapping Communities (ABCD+o2)

### v1

@@ 1 Introduction
Detecting communities is quite a challenging task. In fact, there is no definition of community that researchers and practitioners agree on. Still, it is widely accepted that a community should induce a graph that is denser than the global density of the network [16]. Numerous community detection algorithms have been developed over the years, using various techniques such as optimizing modularity, removing high-betweenness edges, detecting dense subgraphs, and statistical inference. We direct the interested reader to the survey [14] or one of the numerous books on network science [21].
Most community detection algorithms aim to find a partition of the set of nodes, that is, a collection of pairwise disjoint communities with the property that each node belongs to exactly one of them. This is a natural assumption for many scenarios. For example, most of the employees on LinkedIn work for a single employer. On the other hand, users of Instagram can belong to many social groups associated with their workplace, friends, sports, etc. Researchers might be part of many research groups. A large fraction of proteins belong to several protein complexes simultaneously. As a result, many real-world networks are better modelled as a collection of overlapping communities [29] and, moreover, many community detection problems should be tackled by finding overlapping sets of nodes rather than a partition.
@@
To support the development and analysis of overlapping community detection algorithms, a large and diverse catalogue of networks with ground-truth communities is required for testing, tuning and training. Unfortunately, there are few such datasets with communities properly identified and labelled. As a result, there is a need for synthetic random graph models with community structure that resemble real-world networks to benchmark and tune unsupervised clustering algorithms. The popular LFR (Lancichinetti, Fortunato, Radicchi) model [26, 25] generates networks with communities and, at the same time, allows for heterogeneity in the distributions of both node degrees and of community sizes. Due to this structural freedom, the LFR model became a standard and extensively used model for generating artificial networks with ground-truth communities.
@@
A model similar to LFR, the Artificial Benchmark for Community Detection (ABCD[20], was recently introduced and implemented [^1] https://github.com/bkamins/ABCDGraphGenerator.jl/, along with a faster and multithreaded implementation [^2] https://github.com/tolcz/ABCDeGraphGenerator.jl/ (ABCDe[18]. Undirected variants of LFR and ABCD produce graphs with comparable properties, but ABCD (and especially ABCDe) is faster than LFR and can be easily tuned to allow the user to make a smooth transition between the two extremes: pure (disjoint) communities and random graphs with no community structure. Moreover, ABCD is easier to analyze theoretically—for example, in [19] various theoretical asymptotic properties of the are investigated, including the modularity function that, despite some known issues such as the “resolution limit” reported in [15], is an important graph property of networks in the context of community detection. In [5], some interesting and desired self-similar behaviour of the ABCD model is discovered; namely, that the degree distribution of ground-truth communities is asymptotically the same as the degree distribution of the whole graph (appropriately normalized based on their sizes). Finally, the building blocks in the model are flexible and may be adjusted to satisfy different needs. Indeed, the original ABCD model was adjusted to include outliers (ABCD+o[22] and was generalized to hypergraphs (h–ABCD[23] [^3] https://github.com/bkamins/ABCDHypergraphGenerator.jl. For these reasons ABCD is gaining recognition as a benchmark for community detection algorithms. For example, in [2] the authors use both ABCD and LFR graphs to compare 30 community detection algorithms and, in their work, highlight that “being directly comparable to LFR, ABCD offers additional benefits, including higher scalability and better control for adjusting an analogous mixing parameter.”
@@
In this paper we introduce the ABCD+o2 model: a generalization of the ABCD+o model that allows for overlapping communities. The LFR model has been extended in a similar way [25], and in this model the nodes are assigned to communities based on the construction of a random bipartite graph between nodes and communities resulting in (a) a small amount of overlap between almost every pair of communities, and (b) rarely any pair of communities with a large overlap. In ABCD+o2, we instead generate overlapping communities based on a hidden, low-dimensional geometric layer which tends to yield fewer and larger overlaps. This geometric approach to community structure is justified by the fact that latent geometric spaces are believed to shape complex networks (e.g., social media networks shaped by users’ opinions, education, knowledge, interests, etc.). These latent spaces have been successfully employed for many years to model and explain network properties such as self-similarity [32], homophily and aversion [17]. For more details, we direct the reader to the survey [8] or the book [31]. In addition to the geometric layer, the ancillary benefits of the ABCD model (an intuitive noise parameter, a fast implementation, and theoretical analysis) are still present, making the ABCD+o2 model an attractive option for benchmarking community detection algorithms.
@@
2 ABCD+o2 — ABCD with Overlapping Communities and Outliers
@@ 2.1 Notation
For a given $n\in{\mathbb{N}}:=\{1,2,\ldots\}$, we use $[n]$ to denote the set consisting of the first $n$ natural numbers, that is, $[n]:=\{1,2,\ldots,n\}$.
@@
[…]
@@ 2.2 The Configuration Model
The well-known configuration model is an important ingredient of all variants of the ABCD models, so let us formally define it here. Suppose that our goal is to create a graph on $n$ nodes with a given degree distribution $\textbf{d}:=(d_{i},i\in[n])$, where d is a sequence of non-negative integers such that $m:=\sum_{i\in[n]}d_{i}$ is even. We define a random multi-graph $\mathrm{CM}(\textbf{d})$ with a given degree sequence known as the configuration model (sometimes called the pairing model), which was first introduced by Bollobás [9]. (See [6, 35, 34] for related models and results.)
@@ 2.3 Parameters of the ABCD+o2 Model
The following parameters govern the ABCD+o2 model.
$\begin{array}[]{|l|l|l|}\hline\cr\text{Parameter}&\text{Range}&\text{Description}\\ \hline\cr n&{\mathbb{N}}&\text{Number of nodes}\\ s_{0}&{\mathbb{N}}&\text{Number of outliers}\\ \eta&[1,\infty)&\text{Average number of communities a non-outlier node is partof }\\ d&{\mathbb{N}}&\text{Dimension of reference layer}\\ \rho&[-1,1]&\text{Pearson correlation between the degree of nodes and}\\ &&\text{the number of communities they belong to}\\ \hline\cr\gamma&(2,3)&\text{Power-law degree distribution with exponent }\gamma\\ \delta&{\mathbb{N}}&\text{Min degree as least }\delta\\ \Delta&{\mathbb{N}}\setminus[\delta-1]&\text{Max degree at most }\Delta\\ \hline\cr\beta&(1,2)&\text{Power-law community size distribution with exponent}\beta\\ s&{\mathbb{N}}\setminus[\delta]&\text{Min community size at least }s\\ S&{\mathbb{N}}\setminus[s-1]&\text{Max community size at most }S\\ \hline\cr\xi&[0,1]&\text{Level of noise}\\ \hline\cr\end{array}$
@@
Note that the ranges for parameters $\gamma$ and $\beta$ are suggestions chosen according to experimental values commonly observed in complex networks [3, 28]. In fact, users may inject any degree sequence and sequence of community sizes as inputs to the model.
@@ 2.4 Big Picture
The ABCD+o2 model generates a random graph on $n$ nodes with degree sequence $(d_{i},i\in[n])$ and community size sequence $(s_{i},i\in[L])$ following truncated power laws with exponents $\gamma$ and, respectively, $\beta$. There are $s_{0}$ outliers and $\hat{n}=n-s_{0}$ non-outliers. The non-outliers span a family of $L$ communities $(C_{j},j\in[L])$ with each non-outlier belonging to at least one such community. These communities are generated, and will overlap (unless $\eta=1$), according to a hidden $d$-dimensional reference layer in such a way that non-outliers will belong to $\eta$ communities, on average. The non-outliers, with their respective degrees, populate $(C_{j},j\in[L])$ randomly with the caveat that high degree nodes cannot enter small communities unless absolutely necessary. Moreover, one may insist on a correlation between the degree of a node and the number of communities it belongs to. This correlation is controlled by the parameter $\rho$.
Parameter $\xi\in[0,1]$ dictates the amount of noise in the network. Each non-outlier node $i$ has its degree $d_{i}$ split into two parts: community degree $Y_{i}$ and background degree $Z_{i}$ (and thus $d_{i}=Y_{i}+Z_{i}$). Using a random rounding technique, we choose the community degrees and background degrees so that ${\mathbb{E}}\left[Y_{i}\right]=(1-\xi)d_{i}$ and ${\mathbb{E}}\left[Z_{i}\right]=\xi d_{i}$. Note that the neighbours of outliers are sampled from the entire graph, ignoring the underlying community structure, meaning $Y_{i}=0$ and $Z_{i}=d_{i}$ if $i$ is an outlier.
Once nodes are assigned to communities and their degrees are split, the edges of each community are then independently generated by the configuration model on the corresponding community degree sequences, i.e., community $C$ is generated by the configuration on the degree sequence $(Y_{i},i\in C)$. Once communities are generated, the background graph is generated by the configuration model on the degree sequence $(Z_{i},i\in[n])$. The final ABCD+o2 model, after an additional clean-up phase to rewire self-loops and duplicate edges, is the union of the community graphs and the background graph.
@@ 2.5 The ABCD+o2 Construction
This phase is the same as in the original ABCD model and its generalization, ABCD+o. By default, we sample $d_{i}\sim\mathcal{P}\left(\gamma,\delta,\Delta\right)$, independently for each $i\in[n]$, then re-label the samples so that $d_{1}\geq\dots\geq d_{n}$. To ensure that $\sum_{i\in[n]}d_{i}$ is even, we decrease $d_{1}$ by 1 if necessary; we relabel as needed to ensure that $d_{1}\geq\dots\geq d_{n}$. Alternatively, the degree sequence can be given explicitly as an input.
@@
This phase is also the same as in the ABCD+o model. As mentioned in the big picture summary, the neighbours of outliers will be sampled from the entire graph, ignoring the underlying community structure. It feels that this part is straightforward, but a problem might occur when $\xi$ is close to zero. In the extreme case, when $\xi=0$, only the outliers have a non-zero degree in the background graph, and for a simple background graph to exist all outliers must have degree at most $s_{0}-1$. To handle this issue, we restrict which degrees can be assigned to outliers in the following way. We have that $\ell=\sum_{i\in[n]}\min(1,\xi d_{i})$ is a lower bound for the expected number of nodes with positive degree in the background graph. Moreover, the $s_{0}$ outliers will have a positive degree in the background graph. Thus, we expect $\ell+(n-\ell)(s_{0}/n)$ nodes to have positive degree in the background graph, assuming outliers are selected uniformly at random. Therefore, we insist that a node $i$ of degree $d_{i}$ cannot become an outlier unless
@@
By the end of Phase 2, we have a degree sequence $(d_{i},i\in[n])$ and an assignment of degrees to outliers and non-outliers. It is important to keep in mind that, although communities are created in this phase, we do not delegate degrees to these communities until Phase 4. We instead construct a collection of elements which will be points in $d$-dimensional space, each assigned to some number of communities. We will construct $\hat{n}=n-s_{0}$ elements and they will belong to $\eta\geq 1$ communities on average. To be compatible with the original ABCD model, each element will belong to a single primary community, and these primary memberships will partition the elements. Then, we will grow each community by a factor of $\eta$, according to the geometric layer, so that the collective size of all communities is equal (in expectation) to $\eta\hat{n}=\eta(n-s_{0})$. For each element $v$ added to community $C$ during the growth process, $C$ will be referred to as a secondary community of $v$.
Similar to the degree sequence, the community size sequence $(s_{j},j\in[L])$ follows a power-law with parameter $\beta$, minimum value $s$, and maximum value $S$. Hence, the sequence of primary communities $(\hat{s}_{j},j\in[L])$ will need to follow a power-law with parameter $\beta$, minimum value $\hat{s}=\lceil s/\eta\rceil$, and maximum value $\hat{S}=\lfloor S/\eta\rfloor$. In addition, we require $\sum_{j\in[L]}\hat{s}_{j}=\hat{n}$. To satisfy both requirements, we sample $\hat{s}_{j}\sim\mathcal{P}\left(\beta,\hat{s},\hat{S}\right)$ independently until the sum of samples is at least $\hat{n}$. If, at this point, the sum is $\hat{n}+a$ with $a>0$ then we perform one of two actions. If the last added sample has size at least $a+\hat{s}$, we reduce it by $a$. Otherwise (that is, if $c<a+\hat{s}$), we delete this sample, select $c-a$ old samples uniformly at random and increase each by 1. Lastly, let $L$ be the random variable counting the number of communities, and relabel $(\hat{s}_{j},j\in[L])$ so that $\hat{s}_{1}\geq\dots\geq\hat{s}_{L}$. Similar to the degree sequence, the sequence of community sizes can instead be given explicitly as an input. Each primary community of size $\hat{s}_{j}$ will grow to size $s_{j}=\left\lfloor\eta\hat{s}_{j}\right\rceil$ where, for $a\in\mathbb{Z}$ and $b\in[0,1)$, the random variable $\left\lfloor a+b\right\rceil$ is defined as
$\left\lfloor a+b\right\rceil:=\bigg{\{}\begin{array}[]{ll}a&\text{ with probability }1-b,\text{ and}\\ a+1&\text{ with probability }b\,.\end{array}$
Note that ${\mathbb{E}}\left[\left\lfloor a+b\right\rceil\right]=a(1-b)+(a+1)b=a+b$. We get the desired result:
${\mathbb{E}}\left[\sum_{j\in[L]}s_{j}\right]=\sum_{j\in[L]}{\mathbb{E}}\left[s_{j}\right]=\eta\sum_{j\in[L]}\hat{s}_{j}=\eta\hat{n}=\eta(n-s_{0}).$
@@
We now grow each primary community $\hat{C}_{j}$ of size $\hat{s}_{j}$ to form the full community $C_{j}$ of size $s_{j}$. Initially, we set $C_{j}=\hat{C}_{j}$ for all $j\in[L]$. Each community will grow independently, meaning we can grow them in any order (or in parallel). As before, let $R$ be the set of $\hat{n}$ elements. For primary community $\hat{C}_{j}$, let $\mathbf{x}_{j}\in\mathbb{R}^{d}$ be the center of mass (mean) of $\hat{C}_{j}$. We investigate elements of $R$ in order of increasing distance from $\mathbf{x}_{j}$. During the investigation, if an element $v\in R$ is not a member of $\hat{C}_{j}$, then we assign $C_{j}$ as a secondary community of $v$. We stop investigating once the number of members in $C_{j}$ (both primary and secondary) is $s_{j}$. Write $\eta_{v}$ for the number of communities containing element $v\in R$.
@@
At the end of the process, each element is assigned to exactly one primary community and some (possibly zero) secondary communities. In Figure 1 we show an example of the reference layer on $\hat{n}=150$ elements and three communities. Each of the three primary communities in this example consists of 50 elements before growing by a factor of $\eta=2.0$, attracting an additional 50 elements as its secondary members. From a computational perspective, to ensure that finding primary and secondary communities is efficient, we use k-d trees (short for k-dimensional tree) to perform spatial lookups of the sampled points [7].
Figure 1: Example of the reference layer (in $d=2$ dimension) on $\hat{n}=150$ elements consisting of 3 overlapping communities with equal sizes and $\eta=2.0$.
Phase 4: assigning degrees to nodes.
@@
We now have a degree sequence $(d_{i},i\in[n])$, an assignment of degrees to outliers and non-outliers, and a collection of overlapping communities containing elements, each element belonging to one primary community and some number of secondary communities. Let $\hat{\textbf{d}}_{\hat{n}}$ be the subsequence (of length $\hat{n}$) of $(d_{i},i\in[n])$ corresponding to the non-outliers. We are now ready to assign degrees in $\hat{\textbf{d}}_{\hat{n}}$ to the set of elements $R$ which will give us the nodes of the graph. Each element $v\in R$, once paired with node $i$ of degree $d_{i}$, is expected to have $\xi d_{i}$ neighbours in the background graph and the remaining $(1-\xi)d_{i}$ neighbours split evenly between its $\eta_{v}$ communities. Note that, although each element has a distinct primary community, its degree will be split with no preference given to said primary community.
@@
Similar to the potential problem with outliers, we want to avoid large degrees being assigned to small communities. If degree $d_{i}$ is paired with element $v$ then, for each community $C_{j}$ containing $v$, $s_{j}$ must be at least $(1-\xi)d/\eta_{v}+1$ to accommodate for the neighbours of $v$. This is in fact a lower bound as additional neighbours of $v$ coming from the background graph might end up in $C_{j}$ as well. To make additional room in the community graph, a small correction is introduced in both ABCD and ABCD+o guided by the parameter

### v2

@@ 1 Introduction
Detecting communities is quite a challenging task. In fact, there is no definition of community that researchers and practitioners agree on. Still, it is widely accepted that a community should induce a graph that is denser than the global density of the network [girvan2002community]. Numerous community detection algorithms have been developed over the years, using various techniques such as optimizing modularity, removing high-betweenness edges, detecting dense subgraphs, and statistical inference. We direct the interested reader to the survey [fortunato2010community] or one of the numerous books on network science [kaminski2021mining].
Most community detection algorithms aim to find a partition of the set of nodes, that is, a collection of pairwise disjoint communities with the property that each node belongs to exactly one of them. This is a natural assumption for many scenarios. For example, most of the employees on LinkedIn work for a single employer. On the other hand, users of Instagram can belong to many social groups associated with their workplace, friends, sports, etc. Researchers might be part of many research groups. A large fraction of proteins belong to several protein complexes simultaneously. As a result, many real-world networks are better modelled as a collection of overlapping communities [palla2005uncovering] and, moreover, many community detection problems should be tackled by finding overlapping sets of nodes rather than a partition.
@@
To support the development and analysis of overlapping community detection algorithms, a large and diverse catalogue of networks with ground-truth communities is required for testing, tuning and training. Unfortunately, there are few such datasets with communities properly identified and labelled. As a result, there is a need for synthetic random graph models with community structure that resemble real-world networks to benchmark and tune unsupervised clustering algorithms. The popular LFR (Lancichinetti, Fortunato, Radicchi) model [lancichinetti2008benchmark, lancichinetti2009benchmarks] generates networks with communities and, at the same time, allows for heterogeneity in the distributions of both node degrees and of community sizes. Due to this structural freedom, the LFR model became a standard and extensively used model for generating artificial networks with ground-truth communities.
@@
A model similar to LFR, the Artificial Benchmark for Community Detection (ABCD[kaminski2021artificial], was recently introduced and implemented [^1] https://github.com/bkamins/ABCDGraphGenerator.jl/, along with a faster and multithreaded implementation [^2] https://github.com/tolcz/ABCDeGraphGenerator.jl/ (ABCDe[kaminski2022properties]. Undirected variants of LFR and ABCD produce graphs with comparable properties, but ABCD (and especially ABCDe) is faster than LFR and can be easily tuned to allow the user to make a smooth transition between the two extremes: pure (disjoint) communities and random graphs with no community structure. Moreover, ABCD is easier to analyze theoretically—for example, in [kaminski2022modularity] various properties of the ABCD model are investigated, including the modularity which, despite some known issues such as the “resolution limit” reported in [fortunato2007resolution], is an important graph property in the context of community detection. In [self-similarityABCD], some interesting and desired self-similar behaviour of the ABCD model is discovered; namely, that the degree distribution of ground-truth communities is asymptotically the same as the degree distribution of the whole graph (appropriately normalized based on their sizes). Finally, the building blocks in the model are flexible and may be adjusted to satisfy different needs. Indeed, the original ABCD model was adjusted to include outliers (ABCD+o[kaminski2023artificial] and was generalized to hypergraphs (h–ABCD[kaminski2023hypergraph] [^3] https://github.com/bkamins/ABCDHypergraphGenerator.jl. For these reasons ABCD is gaining recognition as a benchmark for community detection algorithms. For example, in [aref2022bayan] the authors use both ABCD and LFR graphs to compare 30 community detection algorithms and, in their work, highlight that “being directly comparable to LFR, ABCD offers additional benefits, including higher scalability and better control for adjusting an analogous mixing parameter.”
@@
In this paper we introduce the ABCD+o2 model: a generalization of the ABCD+o model that allows for overlapping communities. The LFR model has been extended in a similar way [lancichinetti2009benchmarks], and in this model the nodes are assigned to communities based on the construction of a random bipartite graph between nodes and communities resulting in (a) a small amount of overlap between almost every pair of communities, and (b) rarely any pair of communities with a large overlap. In ABCD+o2, we instead generate overlapping communities based on a hidden, low-dimensional geometric layer which tends to yield fewer and larger overlaps. This geometric approach to community structure is justified by the fact that latent geometric spaces are believed to shape complex networks (e.g., social media networks shaped by users’ opinions, education, knowledge, interests, etc.). These latent spaces have been successfully employed for many years to model and explain network properties such as self-similarity [serrano2008self], homophily and aversion [henry2011emergence]. For more details, we direct the reader to the survey [boguna2021network] or the book [serrano2021shortest]. In addition to the geometric layer, the ancillary benefits of the ABCD model (an intuitive noise parameter, a fast implementation, and theoretical analysis) are still present, making the ABCD+o2 model an attractive option for benchmarking community detection algorithms.
As mentioned above, modelling complex networks using underlying geometric hidden layers is not new. However, in most spatial models such as the Spatial Preferential Attachment (SPA) model [aiello2008spatial], nodes are embedded in a metric space, and link formation is influenced by the metric distance between vertices. The metric space is meant to be a feature space, where coordinates represent information associated with the node. For example, in text mining, documents are commonly represented as vectors in a word space. The metric is chosen so that metric distance represents similarity, i.e. nodes whose information entities are closely related will be at a short distance from each other in the metric space. In the ABCD+o2 model, the underlying geometric space plays a slightly different role. In this case, the geometric layer affects the formation of communities which group together nodes that are close to each other. For example, communities in social networks tend to group together users with similar interests, age, geographical locations, all of which can be represented as vectors in a metric space. This natural mechanism creates correlated overlapping clusters which might be challenging to obtain via other means. Connections between users in such social networks are formed as a consequence of such underlying overlapping structure. We are not aware of similar approach used in other models but, given it is quite natural, we would not be surprised if similar techniques were already used.
@@
2 ABCD+o2ABCD with Overlapping Communities and Outliers
@@ 2.1 Notation
For a given $n\in{\mathbb{N}}:=\{1,2,\ldots\}$, we use $[n]$ to denote the set consisting of the first $n$ natural numbers, that is, $n]:=\{1,2,\ldots,n\}$.
@@
For any real number $x=a+b$ with $a\in\mathbb{Z}$ and $b\in[0,1)$, the random variable $\left\lfloor x\right\rceil$ is defined as
$\left\lfloor x\right\rceil:=\bigg\{\begin{array}[]{ll}a&\text{ with probability }1-b,\text{ and}\\ a+1&\text{ with probability }b\,.\end{array}$
Note that ${\mathbb{E}}\left[\left\lfloor x\right\rceil\right]=a(1-b)+(a+1)b=x$.
@@
$\mathbb{P}\left(X=k\right)=\frac{\int_{k}^{k+1}x^{-\gamma}\,dx}{\int_{\delta}^{\Delta+1}x^{-\gamma}\,dx}\,.$ (1)
@@ 2.2 The Configuration Model
The well-known configuration model is an important ingredient of all variants of the ABCD models, so let us formally define it here. Suppose that our goal is to create a graph on $n$ nodes with a given degree sequence $\textbf{d}:=(d_{i},i\in[n])$, where d is a sequence of non-negative integers such that $m:=\sum_{i\in[n]}d_{i}$ is even. We define a random multi-graph $\mathrm{CM}(\textbf{d})$ with a given degree sequence known as the configuration model (sometimes called the pairing model), which was first introduced by Bollobás [bollobas1980probabilistic]. (See [bender1978asymptotic, wormald1984generating, wormald1999models] for related models and results.)
@@ 2.3 Parameters of the ABCD+o2 Model
Table 1 summarizes the parameters that govern the ABCD+o2 model. Note that the ranges for parameters $\gamma$ and $\beta$ are suggestions chosen according to experimental values commonly observed in complex networks [barabasi2016network, orman2009comparison]. In fact, users may inject any degree sequence and sequence of community sizes as inputs to the model.
[…]
@@ 2.4 Big Picture
The ABCD+o2 model generates a random graph on $n$ nodes with a degree sequence and a community size sequence that follow truncated power laws with exponents $\gamma$ and, respectively, $\beta$. The degree sequence is labelled as $(d_{i},i\in[n])$ and the community size sequence is labelled as $(s_{j},j\in[L])$, and note that $L$ is a random variable. The nodes are split into $s_{0}$ outliers and $\hat{n}=n-s_{0}$ non-outliers. The non-outliers span a collection of $L$ communities $(C_{j},j\in[L])$, with $|C_{j}|=s_{j}$ and with each non-outlier belonging to at least one community. These communities are generated, and will overlap (unless $\eta=1$), according to a hidden $d$-dimensional reference layer in such a way that non-outliers will belong to $\eta$ communities, on average. The model allows for a correlation $\rho$ between the degree of a node and the number of communities it belongs to. In some cases, though, a correlation of $\rho$ is unachievable in which case we achieve the highest correlation possible.
Parameter $\xi\in[0,1]$ dictates the amount of noise in the network, i.e., the fraction of edges that do not lie in a single community. Each non-outlier node $i$ has its degree $d_{i}$ split into two parts: community degree $Y_{i}$ and background degree $Z_{i}$ (and thus $d_{i}=Y_{i}+Z_{i}$). Using a random rounding technique, we choose the community degrees and background degrees so that ${\mathbb{E}}\left[Y_{i}\right]=(1-\xi)d_{i}$ and ${\mathbb{E}}\left[Z_{i}\right]=\xi d_{i}$. Note that the neighbours of outliers are sampled from the entire graph, ignoring the underlying community structure, meaning $Y_{i}=0$ and $Z_{i}=d_{i}$ if $i$ is an outlier.
Once nodes are assigned to communities and their degrees are split, the edges of each community graph are independently generated by the configuration model on the corresponding community degree sequences, i.e., community graph $G_{j}$ corresponding to community $C_{j}$ is generated by the configuration on the degree sequence $(Y_{i},i\in C_{j})$. Once communities are generated, the background graph is generated by the configuration model on the degree sequence $(Z_{i},i\in[n])$. The final ABCD+o2 model, after an additional clean-up phase to rewire self-loops and duplicate edges, is the union of the community graphs and the background graph.
@@ 2.5 The ABCD+o2 Construction
This phase is the same as in the original ABCD model and its generalization, ABCD+o. By default, we sample $d_{i}\sim\mathcal{P}\left(\gamma,\delta,\Delta\right)$, independently for each $i\in[n]$, then re-label the samples so that $d_{1}\geq\dots\geq d_{n}$. To ensure that $\sum_{i\in[n]}d_{i}$ is even, we decrease $d_{1}$ by 1 if necessary; we relabel again if needed to ensure that the sequence remains monotone. Alternatively, the degree sequence can be given explicitly as an input.
@@
This phase is also the same as in the ABCD+o model. As mentioned in the big picture summary, the neighbours of outliers will be sampled from the entire graph, ignoring the underlying community structure. It feels that this part is straightforward, but a problem might occur when $\xi$ is close to zero. In the extreme case, when $\xi=0$, the edges in the background graph are precisely the edges connecting two outliers, and for a simple background graph to exist all outliers must have degree at most $s_{0}-1$. To handle the more general issue that arises when $\xi$ is small, we need to estimate the number of nodes with at least one edge in the background graph, then use this estimation to upper-bound the degree of an outlier.
The background degree of a node is the number of edges in the background graph containing said node. The expected background degree of node $i$ of degree $d_{i}$ is $\left\lfloor\xi d_{i}\right\rceil$ (recall the definition of $\left\lfloor\cdot\right\rceil$ given in Section 2.1; more details will be provided in Phase 5). In particular, the background degree of node $i$ is positive with probability $\min(1,\xi d_{i})$. Hence, before it is decided which nodes become outliers, we expect $\ell=\sum_{i\in[n]}\min(1,\xi d_{i})$ nodes with positive background degree. Moreover, the $s_{0}$ outliers are guaranteed to have positive background degree. Thus, in total we expect $\ell+(n-\ell)(s_{0}/n)$ nodes with positive background degree. Therefore, we insist that a node $i$ of degree $d_{i}$ cannot become an outlier unless
@@
By the end of Phase 2, we have a degree sequence $(d_{i},i\in[n])$ and an assignment of degrees to outliers and non-outliers. It is important to keep in mind that, although communities are created in this phase, we do not delegate degrees to these communities until Phase 4. We instead construct a collection of elements which will be points in $d$-dimensional space, each assigned to some number of communities. We will construct $\hat{n}=n-s_{0}$ elements and they will belong to $\eta\geq 1$ communities on average. To be compatible with the original ABCD model, each element will belong to a single primary community, and these primary memberships will partition the elements. Then, we will grow each community by a factor of $\eta$, adding nearby elements in the $d$-dimensional space, so that the collective size of all communities is equal (in expectation) to $\eta\hat{n}=\eta(n-s_{0})$ (we define this growth process formally below). For each element $v$ added to community $C$ during the growth process, $C$ will be referred to as a secondary community of $v$.
@@
${\mathbb{E}}\left[\sum_{j\in[L]}s_{j}\right]=\sum_{j\in[L]}{\mathbb{E}}\left[s_{j}\right]=\sum_{j\in[L]}{\mathbb{E}}\left[\left\lfloor\eta\hat{s}_{j}\right\rceil\right]=\eta\sum_{j\in[L]}\hat{s}_{j}=\eta\hat{n}=\eta(n-s_{0}),$
as desired.
We now create our $d$-dimensional reference layer that will guide the process of assigning elements to communities. One may think of this reference layer as a map representing various latent properties of elements (such as people’s age, education, geographic location, beliefs, etc.) shaping communities (such as communities in social media) based on proximity. In this reference layer, each of the $\hat{n}$ elements is generated as a vector in $\mathbb{R}^{d}$, sampled independently and uniformly at random from the ball of radius $1$ centred at the origin $\textbf{0}=(0,0,\ldots,0)$. Let $R$ be the set of $\hat{n}$ elements and initalize $R_{1}=R$. We assign elements to their primary communities, dealing with one primary community at a time. To form primary community $\hat{C}_{j}$, we first select the element in $R_{j}$ furthest from 0. This element, together with its $\hat{s}_{j}-1$ nearest neighbours in $R_{j}$, form the elements of $\hat{C}_{j}$. We then remove these elements from $R_{j}$ to construct $R_{j+1}$ and repeat the process until all elements have been assigned a primary community. At the end of the process, each element belongs to exactly one primary community and the collection $(\hat{C}_{j},j\in[L])$ partitions the elements.
@@
We are now ready to define the growth process as mentioned above. We grow each primary community $\hat{C}_{j}$ of size $\hat{s}_{j}$ to form the full community $C_{j}$ of size $s_{j}$. Initially, we set $C_{j}=\hat{C}_{j}$ for all $j\in[L]$. Each community will grow independently, meaning we can grow them in any order (or in parallel). As before, let $R$ be the set of $\hat{n}$ elements. For primary community $\hat{C}_{j}$, let $\mathbf{x}_{j}\in\mathbb{R}^{d}$ be the center of mass (mean) of $\hat{C}_{j}$. We investigate elements of $R$ in order of increasing distance from $\mathbf{x}_{j}$. During the investigation, if an element $v\in R$ is not a member of $\hat{C}_{j}$, then we assign $C_{j}$ as a secondary community of $v$. We stop investigating once the number of members in $C_{j}$ (both primary and secondary) is $s_{j}$. Write $\eta_{v}$ for the number of communities containing element $v\in R$.
@@
At the end of the process, each element is assigned to exactly one primary community and some (possibly zero) secondary communities. In Figure 1 we show an example of the reference layer on $\hat{n}=150$ elements and three communities. Each of the three primary communities in this example consists of 50 elements before growing by a factor of $\eta=2.0$, attracting an additional 50 elements as its secondary members. From a computational perspective, to ensure that finding primary and secondary communities is efficient, we use k-d trees (short for k-dimensional tree) to perform spatial lookups of the sampled points [bentley1975kdtrees].
Figure 1: Example of assigning the community $C_{3}$ (Primary 2 in the legend) with $\hat{s}_{3}=40$ and $\eta=1.75$ using a 2 dimensional reference layer. First, we select the element furthest from the origin that does not have a primary community as the seed. Next, we assign $C_{3}$ as the primary community of the seed, as well as the $\hat{s}_{3}-1$ nearest neighbours of the seed that do not yet have a primary community. Finally, we expand $C_{3}$ by a factor of $\eta$ (from $40$ to $40\cdot 1.75=70$ nodes) by taking the nearest neighbours to the primary center of mass.
Phase 4: assigning degrees to elements.
@@
We now have a degree sequence $(d_{i},i\in[n])$, an assignment of degrees to outliers and non-outliers, and a collection of overlapping communities containing elements, each element belonging to one primary community and some number of secondary communities. Let $\hat{\textbf{d}}_{\hat{n}}$ be the subsequence (of length $\hat{n}$) of $(d_{i},i\in[n])$ corresponding to the non-outliers. We are now ready to assign degrees in $\hat{\textbf{d}}_{\hat{n}}$ to the set of elements $R$ which will give us the non-outlier nodes of the graph.
@@
Similar to the potential problem with outliers, we want to avoid large degrees being assigned to small communities. When we create edges in Phase 5, element $v$, paired with some degree $d_{i}$, will have at least $\lfloor\xi d_{i}\rfloor$ of its degree delegated to the background graph, and will split the remainder of its degree evenly among its $\eta_{v}$ communities. Thus, if degree $d_{i}$ is paired with element $v$ then, for each community $C_{j}$ containing $v$, $|C_{j}|=s_{j}$ must be at least $\Big\lceil\lceil(1-\xi)d_{i}\rceil/\eta_{v}\Big\rceil+1$ to accommodate the neighbours of $v$. This is in fact a lower bound as additional neighbours of $v$ coming from the background graph might end up in $C_{j}$ as well. This issue was addressed in the original ABCD paper [kaminski2021artificial] where, to make room for the extra neighbours, a small correction was introduced, guided by the parameter